OpenStudy (anonymous):

Hi, can someone please explain how they would derive f(x)=x^(2/3)*(x^2-4)

OpenStudy (freckles):

you want to differentiate? if so, use product rule first.

OpenStudy (jhannybean):

derive or derivate? your question is kind of vague!

OpenStudy (jhannybean):

Omg, everything I typed disappeared because I pressed the back button -.-

OpenStudy (freckles):

hate when that happens

OpenStudy (jhannybean):

\[f(x)=x^{2/3}(x^2-4)\]your \(f = x^{2/3}\) and your \(g=(x^2-4)\) so yes, you would use product rule as @freckles stated :)

OpenStudy (anonymous):

\[\frac{ d }{ dx }(x^\frac{ 2 }{ 3 })(x^2-4) = \frac{ 2 }{ 3 }x^{(-\frac{ 1 }{ 3 })}(x^2-4)+x ^{2/3}(2x)\]

OpenStudy (anonymous):

It just got so complicated, and I got lost at this part. :(

OpenStudy (freckles):

\[(fg)'=f'g+fg' \\ [(x^\frac{2}{3})(x^2-4)]'=(x^\frac{2}{3})'(x^2-4)+x^\frac{2}{3}(x^2-4)'\]

OpenStudy (freckles):

so first what are \[(x^\frac{2}{3})' =? \\ (x^2-4)'=?\]

OpenStudy (anonymous):

2x?

OpenStudy (freckles):

oh it loks like you already have the derivatives above

OpenStudy (freckles):

looks*

OpenStudy (anonymous):

Yes I just dont know how to make it simpler

OpenStudy (anonymous):

Im looking for the minimum and maximum values

OpenStudy (freckles):

\[\frac{ d }{ dx }(x^\frac{ 2 }{ 3 })(x^2-4) = \frac{ 2 }{ 3 }x^{(-\frac{ 1 }{ 3 })}(x^2-4)+x ^{2/3}(2x) \\ x^\frac{-1}{3} [ \frac{2}{3}(x^2-4)+x^\frac{3}{3}2x ]\]

OpenStudy (freckles):

that should help

OpenStudy (jhannybean):

Oh, then you want to change your function the have all positive values. So then, you can solve for your critical points and test them

OpenStudy (jhannybean):

Your local mins and max's can be found by the first derivative test, but first you need to find your crit points.

OpenStudy (anonymous):

^ I was hoping to get that answer.

OpenStudy (jhannybean):

mmhmm. letme resolve it so you can see how it works.

OpenStudy (jhannybean):

\[ f(x)=x^{2/3}(x^2-4)\]\[f'(x) = \frac{2}{3}x^{-1/3}(x^2-4) +2x(x^{2/3})\]\[f'(x) = \frac{2(x^2-4)}{3x^{1/3}} +2x^{5/3}\]\[f'(x) = \frac{2(x^2-4)}{3x^{1/3}}+\frac{2x^{5/3}\cdot 3x^{1/3}}{3x^{1/3}}\]\[f'(x) = \frac{2x^2-8+6x^2}{3x^{1/3}}\]\[f'(x) = \frac{8x^2-8}{3x^{1/3}}\]\[f'(x)=\frac{8(x^2-1)}{3x^{1/3}}\]

OpenStudy (jhannybean):

You just kind of have to keep going with it.

OpenStudy (jhannybean):

Oh and \(x^{1/3} = \sqrt[3]{x}\)

OpenStudy (anonymous):

Aaaaamzing!!!! Thank you!!!

OpenStudy (anonymous):

Woooot!

OpenStudy (jhannybean):

Can you find your critical points and then use them to find your ma and mins? :)

OpenStudy (anonymous):

Im just looking between [-1,2] so if x=1 then there the tangent line slop at that point is 0. At 0.5 the slop is -2.5 and at 2 the slop is 6.3, so x=1 must be the minimum, and x=2 must be the maximum. When I plug these x values back into the original equation I get f(1)=-3 and f(2)=0

OpenStudy (anonymous):

But what about f(0)? That would also equal to a minimum but f'(0) D.N.E....

OpenStudy (freckles):

you should have 3 critical numbers

OpenStudy (freckles):

what values did you get for critical numbers?

OpenStudy (freckles):

you find when f'=0 and when f' dne the values that satisfy those equations where the x values actually exist for the original function are your critical numbers

OpenStudy (freckles):

\[\frac{ d }{ dx }(x^\frac{ 2 }{ 3 })(x^2-4) = \frac{ 2 }{ 3 }x^{(-\frac{ 1 }{ 3 })}(x^2-4)+x ^{2/3}(2x) \\ x^\frac{-1}{3} [ \frac{2}{3}(x^2-4)+x^\frac{3}{3}2x ] \\ \frac{1}{x^\frac{1}{3}} (\frac{2}{3}x^2-\frac{8}{3}+2x^2) \\ \frac{2}{x^\frac{1}{3}} (\frac{1}{3}x^2-\frac{4}{3}+x^2) \\ \frac{2}{x^\frac{1}{3}}(\frac{4}{3}x^2-\frac{4}{3}) \\ \frac{8}{3 x^\frac{1}{3}}(x^2-1)\] so x^2-1=0 when x=? or x=? x^(1/3)=0 when x=?

OpenStudy (freckles):

and yes I know @Jhannybean already put it into that form I was just finishing the way I started earlier :)

OpenStudy (jhannybean):

:P Of course. Both methods work!

OpenStudy (jhannybean):

\[f'(x)=\frac{8(x^2-1)}{3x^{1/3}}\]\[8(x^2-1)=0\]\[3x^{1/3}=0\]These should give you two values, 2 crit points from the first function, and the same for the bottom.

OpenStudy (anonymous):

Yes so x=1 and x=-1 are the minima? and x=2 is the maximum?

OpenStudy (jhannybean):

Can you tell me where the x=2 came from?

OpenStudy (freckles):

ok but x^(1/3)=0 when x=?

OpenStudy (anonymous):

Im solving for the domain section of [-1,2]

OpenStudy (freckles):

what number raised to any power (not including 0) will give you 0?

OpenStudy (anonymous):

To my knowledge only zero....

OpenStudy (freckles):

so critical numbers are -1,1,0 endpoints are -1,2 --- now only local extrema can happen at non-endpoint numbers

OpenStudy (jhannybean):

Good, and if you are evaluating at those end points then you would not only test your critical points but those end points as well to mind your maximum/minimum values, which I believe you did?

OpenStudy (anonymous):

Yes, i tested for -1 and 2. 2 was a maximum endpoint

OpenStudy (freckles):

doesn't f(2)=f(0)?

OpenStudy (anonymous):

This is interesting.

OpenStudy (anonymous):

Is it right to say \[x^\frac{ 2 }{ 3 } = x^\frac{ 1 }{ 3 }\times x^ \frac{ 1 }{ 3 }\]

OpenStudy (freckles):

\[\text{ yes you can say that } \text{ or } (x^\frac{1}{3})^2 \]

OpenStudy (jhannybean):

Mmhmm. Because when functions multiply, powers add.

OpenStudy (freckles):

I just want to say you have an absolute max value occurring at two numbers not just one

OpenStudy (anonymous):

So at x= 0, 2?

OpenStudy (freckles):

yes since f(0)=f(2)=0

OpenStudy (anonymous):

The interesting thin is the screenshot i posted. Its from wolfram alpha -1^0.6666 = -1

OpenStudy (anonymous):

But \[-1^{\frac{ 2 }{ 3 }}= 1\]

OpenStudy (jhannybean):

Because you're squaring (-1) and then taking it's cube root

OpenStudy (jhannybean):

Which is simply \(1\)

OpenStudy (freckles):

\[(-1)^\frac{2}{3}=(-1)^2=1 \\ -1^\frac{2}{3}=-1\]

OpenStudy (anonymous):

Yes, I understand.

OpenStudy (freckles):

and wolfram's graph I have seen doesn't always agree with the graphs we make

OpenStudy (freckles):

if that makes sense

OpenStudy (anonymous):

Math is fun like that haha. I would like to think we are right and the computer at wolfram is wrong :D

OpenStudy (anonymous):

But I'm guessing there is more too it

OpenStudy (freckles):